can u please help me with both of these, thanks
Let g(x) = x3 + x. Compute g(x+h)-g(x).( h
Use the given information to compute sin Зп and π<θ < 2 cos 0 = 8
Answer #1

soin :- given, gin) = n3th ginth)- gin) = ? h To find ginth) simply supositi substitute n=nth in the function gen) Thus, glutNow, g(x+h)-gR) MC 3+2 4 ²h+nh²+ nth tn +2442 th3th - (3 + x) h gluth)-gin) + 2n²h tun? +4²n +x+ 2062 th3th x ginth)-ging = 2soint given, coso = ㅎ TO<3 2 since coso=1-asin 20/2 asin 20/2= I-coso sin 2012 = In cosa 2_ Now, 1 sin ²0/2 = 1 - 1 - / 8 2

PLEASE UPVOTE.